22west:

Subtract -b+3 from -11b-4

22west:

@dude

dude:

If you are subtracting you have to rewrite like \(11b-4-(-b+3)\) Distribute the negative first

22west:

So the negative turns into b?

22west:

Positive b?

dude:

What is negative times a negative?

22west:

Negative?

dude:

No, its a positive

22west:

-11b - 4 = 4?

22west:

-15 sorry

22west:

Is that the answer?

dude:

Not quite, we have to distribute first you're going too fast

22west:

Okay

22west:

We only have 7 minutes

dude:



dude:

I cannot help you if this is a form of school assessment.

22west:

No

22west:

Just a question my teacher gave us

dude:

Alright so do \(-1\cdot -11=\) and \(-1\cdot 3=\)

22west:

11 and -3

dude:

Not quite, and also my mistake for wrong numbers Supposed to be \(-1\cdot -1=\) and \(-1\cdot 3=\) Look out for the signs!

22west:

1 and -3

dude:

Right so it should be \(11b-4+1b-3\) Now we combine like terms

22west:

Without the variable b it would be 5

22west:

10 seconds

22west:

Nvm 5 more minutes

dude:



22west:

13b and 1

dude:

Check your math What is 11 + 1? x'D

22west:

12b and 1

22west:

Lol

dude:

-4-3 is not 1

22west:

Sorry only have 2 minutes

22west:

-1

dude:

No, if you have 2 numbers with a negative just add them but keep the negative sign

22west:

Yeah -1

dude:

Add them lol What is 4+3?

22west:

7

dude:

But keep the negative, so -7

22west:

-7

22west:

12b and -7?

dude:

Right so your answer is 12b-7

22west:

Thank you

22west:

Gtg byeeeee

dude:

Alright cya

22west:

@dude

22west:

1 attachment
22west:

I got it wrong

22west:

I guess it was:

1 attachment
dude:

Ah that was my mistake! I forgot about the sign at the beginning \(\color{red}{11b-4+1b-3}\) What I had \(\color{purple}{(-)}\color{green}{11b-4+1b-3}\) i missed the negative

22west:

Okay

22west:

It's fine dude